OpenStudy (anonymous):

(1/(x+h)^2-1/x^2)/h, h cannot equal 0. Please help me find the equivalency

OpenStudy (freckles):

Try combining top fractions

OpenStudy (anonymous):

GIVE ME best response first

OpenStudy (anonymous):

Do you know how to add/subtract fractions?

OpenStudy (aum):

\[ \frac{1}{(x+h)^2} - \frac{1}{x^2} = \frac{x^2-(x+h)^2}{x^2(x+h)^2} = \frac{(x+x+h)(x-x-h)}{x^2(x+h)^2} = \frac{(2x+h)(-h)}{x^2(x+h)^2} \]Divide that by h.

OpenStudy (anonymous):

Thank you so much @aum !!!!!!!!

OpenStudy (anonymous):

\[- \frac{ 2x+h }{ x ^{2}(x+h)^{2} } \]

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

Did you get that? @morganheist

OpenStudy (anonymous):

yes @Jesstho.-. ! Thank you bothhh. It's greatly appreciated

OpenStudy (anonymous):

You're welcome